What is the weight percent (%) acetic acid (CH3CO2H) for a 5.88 m aqueous acetic acid solution?
Answer #1

Consider 1 Kg of solvent

use:

molality = number of mol of CH3CO2H / mass of solvent in Kg

5.88 = number of mol of CH3CO2H / 1 Kg

number of mol of CH3CO2H = 5.88 mol

Molar mass of CH3CO2H,

MM = 2*MM(C) + 4*MM(H) + 2*MM(O)

= 2*12.01 + 4*1.008 + 2*16.0

= 60.052 g/mol

use:

mass of CH3CO2H,

m = number of mol * molar mass

= 5.88 mol * 60.05 g/mol

= 353.1 g

mass % = mass of CH3CO2H * 100 / mass of solution

= 353.1 * 100 / (353.1 + 1000)

= 26.1 %

Answer: 26.1 %